determine the number of de broglie waves made by an electron in1s in third orbit of h atom

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

mvr=nh/2π,n=3,mv2/r=(1/4πε0)e2/r2, de Broglie wavelength λ=h/mv

Solve the above equations to obtain the answer